OpenStudy (anonymous):

help i need make my own pool just like this but with different numbersStep 1: The Family Pool Volume of Pool V = lwh V = 25 • 10 • 8 V = 2000 ft3 Amount of Pool Liner Area of Base + Area of Left Side + Area of Right Side + Area of Front Side + Area of Back Side (25 • 10) + (10 • 8) + (10 • 8) + (25 • 8) + (25 • 8) 250 + 80 + 80 + 200 + 200 810 ft2 Volume of Water V = lwh V = 25 • 10 • 7.5 V = 1875 ft3

OpenStudy (anonymous):

@bngaure @bobbylynn @RioAnne @kmarie1123

OpenStudy (anonymous):

@RioAnne

OpenStudy (anonymous):

@sasogeek

OpenStudy (anonymous):

@bobbylynn

OpenStudy (sasogeek):

ok so basically you know what a pool looks like... right? draw a pool, put numbers that make sense on the dimensions, and solve the rest :) sounds easy :P i'll do an example, then you follow what i do, and lemme know what you get :)

OpenStudy (anonymous):

lol yea thnks

OpenStudy (anonymous):

@Agent_Sniffles

OpenStudy (anonymous):

@yummydum

OpenStudy (sasogeek):

 ok so basically what i did was add 5 to all the sides that was in the example, it's not necessary, it just has to make sense. now if we want the volume of the pool i have here (mind you, that pool will be in my mansion in the future ;) ) all we have to do is find the l * w * h xD hence, volume of saso's beautiful pool = l*w*h =30*15*13 =5850 ft^3 Amount of pool liner = area of left and right side + area of front and back side + area of base area of left side=area of right , area of front side = area of back side ==> Amount of pool liner=2(area of left) + 2(area of front) + area of base =2(13*30) + 2(15*13) + (15*30) =780 + 390 + 450 =1620 ft^2 i'm not sure how you're going to measure the volume of water though because it's my pool and i can vary the height of the water any time i want :/ but that's an example :)